OpenStudy (anonymous):

4.Write the equation of the line which passes through (2, –3) and is perpendicular to y = 4x + 7 in standard form.

OpenStudy (anonymous):

perpendicular slopes are negative reciprocal of each other ( -1 divide by the slope) find the new slope and put it in y = mx + b then substitute (2,-3) ~

OpenStudy (anonymous):

it is perpendicular to y=4x+7.its slope will be -1/4 it also passes through (2,3) so equation is y-3=-1/4(x-2)

Parth (parthkohli):

For the next step, you substitute x with 2 and y with 3.

OpenStudy (anonymous):

y = mx + b -3 = (-1/4)(2) + b -3 = (-1/2) + b -5/2 = b one equation of the line is y = (-1/4)x - 5/2. The standard equation of a line is of the form Ax + By = C. We'll need to do some rearranging: y = (-1/4)x - 5/2 y + (1/4)x = -5/2 4y + x = -10